OpenStudy (alexh107):

Difference quotient for f(x) = x^2+x

zepdrix (zepdrix):

Hey Alex! Do you remember your slope formula?\[\large\rm m=\frac{y_2-y_1}{x_2-x_1}\]Well that's all our difference quotient really is, but it's written with function notation so it looks a little fancier.

zepdrix (zepdrix):

\(\large\rm y=x^2+x\) So with the old notation, we might label our y2 and y1 like this, \(\large\rm y_1=(x_1)^2+x_1\) \(\large\rm y_2=(x_2)^2+x_2\) With function notation we would write the first point like this,\[\large\rm f(x_1)=(x_1)^2+x_1\]and instead of calling the second point x2, it's going to be x1 plus some unknown distance to the right. So our x2 is actually located at x1+h. So our second point will look like this,\[\large\rm f(x_1+h)=(x_1+h)^2+x_1\]

zepdrix (zepdrix):

So then subtracting our y's: \(\large\rm f(x_1+h)-f(x)\) and subtracting the x's below: \(\large\rm x_1+h-x_1\) (Which simplifies to h), gives us our difference quotient.

zepdrix (zepdrix):

Sorry maybe that's more information than you needed :)\[\large\rm \frac{f(x+h)-f(x)}{h}\]

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{x})=(\color{orangered}{x})^2+\color{orangered}{x}\]If we evaluate this function at x+h instead of x,\[\large\rm f(\color{orangered}{x+h})=(\color{orangered}{x+h})^2+\color{orangered}{x+h}\]We can see the things that we'll be plugging into the numerator.

OpenStudy (alexh107):

Okay I'm with you so far as to plugging the x+h in to get f(x+h) = (x+h)^2 + x+h but after that I'm a little lost.

zepdrix (zepdrix):

So we've determined that, \(\large\rm \color{indianred}{f(x)=x^2+x}\) and \(\large\rm \color{royalblue}{f(x+h)=(x+h)^2+(x+h)}\) So we're going to make the replacement in our difference quotient,\[\large\rm \frac{\color{royalblue}{f(x+h)}-\color{orangered}{f(x)}}{h}\quad=\quad \frac{\color{royalblue}{(x+h)^2+(x+h)}-\color{orangered}{(x^2+x)}}{h}\]

zepdrix (zepdrix):

Do you remember how to expand out a square binomial? \(\large\rm (x+h)^2=(x+h)(x+h)=?\)

OpenStudy (alexh107):

I think so, that would be x^2 + 2xh + h^2?

zepdrix (zepdrix):

\[\large\rm \frac{(x+h)^2+(x+h)-(x^2+x)}{h}\]Ok good job, we've expanded the square,\[\large\rm \frac{x^2+2xh+h^2+(x+h)-(x^2+x)}{h}\]Let's drop the brackets on the middle term, those aren't necessary. And let's distribute the negative to each of the last two terms,\[\large\rm \frac{x^2+2xh+h^2+x+h-x^2-x}{h}\]

zepdrix (zepdrix):

From here, combine like-terms.

zepdrix (zepdrix):

If you do it correctly, you'll be left with ONLY h stuff in the numerator.

OpenStudy (alexh107):

Ok one sec

OpenStudy (alexh107):

h^2+2xh+h / h?

zepdrix (zepdrix):

Good good good. \[\large\rm \frac{h^2+2xh+h}{h}\]From here, you can either directly divide h out of everything, or, if you'd like you can factor an h out of each term in the numerator before doing the division.

OpenStudy (alexh107):

Would I get h + 2x?

zepdrix (zepdrix):

If we factor an h out of each term in the numerator, we have,\[\large\rm \frac{h(h+2x+1)}{h}\]Don't forget about that 1 on the end! :) Factoring h out h doesn't leave you with nothing.

zepdrix (zepdrix):

h out of h*

OpenStudy (alexh107):

My bad, is it 2x + 1 then?

zepdrix (zepdrix):

Dividing the h's,\[\large\rm \frac{\cancel h(h+2x+1)}{\cancel h}\quad=\quad h+2x+1\]Like that, ya?

OpenStudy (alexh107):

Oh yeah you're right.

zepdrix (zepdrix):

Yayyy team we did it! It's not so bad, right? Just a few tricky Algebra steps.

OpenStudy (alexh107):

Yeah, you made it a lot easier to understand. Thank you so much!